# Towards Robust Real-World Multivariate Time Series Forecasting: A Unified Framework for Dependency, Asynchrony, and Missingness

### v1

@@ C.2 Comparison of Attention Masking Strategies
Table 10: Comparison of attention masking strategies on the SolarWind dataset under the block-wise test-time missing scenario. Results are reported in CMSE and CMAE, averaged over all prediction lengths with a missing ratio $m=0.125$. Among the methods, CD-ReadOnly and CD-ReadOnly-Indexed achieve the best performance under our proposed practical setting.
Strategies CI-ReadOnly CI-Mutual CD-ReadOnly CD-Mutual CD-ReadOnly-Indexed CD-Mutual-Indexed
CMSE 0.412 0.416 0.396 0.398 0.396 0.399
CMAE 0.464 0.468 0.458 0.458 0.458 0.460
@@ Appendix D Forecasting Results in a Conventional Multivariate Setting
To demonstrate that our model remains effective even under standard multivariate scenarios, we additionally evaluate it on regularly sampled multivariate time series without any missing values. By aligning the sampling period and patch length across all channels, our model can be directly applied to this conventional setting. This evaluation shows that our model remains applicable beyond the proposed practical settings. For comparison, baseline results [4, 42, 39, 29, 15, 23, 34, 53, 7, 44, 50] are obtained either from the original publications or by reproducing them using the TSLib codebase [41]. We compare forecasting performance on five widely used benchmark datasets [55, 43]: ETTh1, ETTh2, ETTm1, ETTm2, and Weather. As shown in Table 11, our model achieves performance comparable to or better than state-of-the-art baselines across all benchmark datasets in the conventional forecasting setting.
@@ Appendix E Full Results of Our Proposed Practical Setting
Table 13 shows the results for Scenario 1 using linearly interpolated datasets, where practical data with channel-wise asynchronous sampling are converted to a regular sampling grid to accommodate baseline models. Table 14 presents the results for Scenario 2, which adds block-wise test-time missing inputs on top of asynchronous sampling. All results include error bars based on 5 random seeds to reflect performance variability. Across both scenarios, our model consistently demonstrates stronger performance than baseline methods.
Table 12: Full results of the Asynchronous long-term multivariate forecasting task (forward fill).
Strategy
Channel-Dependent
Channel-Independent
Models
CTF(ours)
TimeXer
TimeMixer
iTrans.
CrossGNN
TimesNet
PatchTST
DLinear
Metric
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
CMSE
CMAE
ETT1
192
0.336
0.371
0.358
0.386
0.348
0.377
0.355
0.386
0.363
0.382
0.371
0.397
0.344
0.376
0.354
0.378
336
0.363
0.390
0.392
0.412
0.379
0.395
0.394
0.411
0.398
0.402
0.418
0.423
0.376
0.398
0.388
0.397
768
0.427
0.430
0.440
0.440
0.449
0.435
0.463
0.450
0.460
0.439
0.464
0.458
0.447
0.439
0.448
0.436
1152
0.467
0.454
0.479
0.462
0.488
0.461
0.505
0.476
0.502
0.464
0.488
0.466
0.476
0.457
0.494
0.468
Avg
0.398
0.411
0.417
0.425
0.416
0.417
0.429
0.431
0.431
0.421
0.435
0.436
0.411
0.418
0.421
0.420
SolarWind
288
0.322
0.383
0.340
0.415
0.369
0.434
0.368
0.423
0.348
0.418
0.347
0.397
0.332
0.406
0.337
0.442
576
0.405
0.460
0.428
0.483
0.447
0.480
0.451
0.477
0.443
0.473
0.534
0.504
0.410
0.466
0.415
0.512
864
0.417
0.461
0.424
0.465
0.485
0.506
0.489
0.501
0.511
0.512
0.474
0.489
0.448
0.491
0.456
0.545
1152
0.456
0.488
0.489
0.514
0.491
0.522
0.544
0.526
0.572
0.547
0.553
0.525
0.461
0.490
0.469
0.557
Avg
0.400
0.448
0.420
0.469
0.448
0.486
0.463
0.482
0.468
0.488
0.477
0.479
0.413
0.463
0.419
0.514
Weather
144
0.179
0.225
0.222
0.257
0.204
0.243
0.235
0.271
0.196
0.240
0.253
0.285
0.200
0.237
0.207
0.261
288
0.242
0.275
0.261
0.285
0.277
0.296
0.288
0.307
0.265
0.311
0.261
0.288
0.292
0.320
0.260
0.308
576
0.311
0.322
0.331
0.331
0.348
0.346
0.349
0.343
0.326
0.339
0.335
0.332
0.351
0.355
0.320
0.356
864
0.343
0.347
0.361
0.357
0.352
0.355
0.380
0.371
0.366
0.368
0.377
0.366
0.372
0.359
0.355
0.395
Avg
0.269
0.293
0.294
0.308
0.295
0.310
0.313
0.323
0.289
0.314
0.307
0.318
0.304
0.318
0.286
0.330
CHS
120
0.104
0.062
0.129
0.070
0.111
0.062
0.125
0.074
0.106
0.075
0.137
0.074
0.135
0.066
0.111
0.097
240
0.165
0.087
0.177
0.089
0.177
0.089
0.184
0.098
0.168
0.099
0.190
0.099
0.216
0.094
0.186
0.171
480
0.312
0.140
0.321
0.142
0.318
0.139
0.329
0.144
0.310
0.159
0.340
0.148
0.335
0.142
0.375
0.288
840
0.548
0.213
0.559
0.214
0.572
0.216
0.570
0.216
0.549
0.224
0.602
0.224
0.614
0.230
0.695
0.435
Avg
0.282
0.126
0.296
0.129
0.294
0.127
0.302
0.133
0.283
0.139
0.317
0.136
0.325
0.133
0.342
0.248

### v2

@@ C.2 Comparison of Attention Masking Strategies
Table 11: Comparison of attention masking strategies on the SolarWind dataset under the block-wise test-time missing scenario. Results are reported in CMSE and CMAE, averaged over all prediction lengths with a missing ratio $m=0.375$. Among the methods, CD-ReadOnly and CD-ReadOnly-Indexed achieve the best performance under our proposed practical setting.
Strategy CI-ReadOnly CI-Mutual CD-ReadOnly CD-Mutual CD-ReadOnly-Indexed CD-Mutual-Indexed
CMSE 0.474 0.457 0.452 0.456 0.452 0.457
CMAE 0.521 0.512 0.508 0.510 0.508 0.512
@@
C.3 Scalability Analysis
We further investigate the scalability of ChannelTokenFormer (CTF) with respect to both the number of channels and the input sequence length.
Channel scalability.
We measured runtime per training iteration and maximum memory usage as the number of input channels increases. Results in Table 12 show that CTF scales reliably up to 275 channels on a single 24GB GPU. Out-of-memory (OOM) is observed only at 280 channels. Training runtime grows proportionally with channel count (e.g., 2.69s/iter at $C=200$), while inference latency remains comparable to TimeXer (0.903s vs. 0.917s).
Table 12: Channel scalability of CTF on a single GPU with 24GB VRAM. We report training runtime (s/iter) and peak memory usage (MB) as the number of input channels increases. CTF scales reliably up to 275 channels, with out-of-memory (OOM) occurring at 280 channels.
Metric / # of Channels 5 30 50 100 200 275 280
Runtime (s/iter) 0.0186 0.1193 0.2307 0.7173 2.6912 5.4175 OOM
Memory (MB) 108.9 460.3 911.2 2843.9 10386.1 17834.9
Input length scalability.
We also tested longer input horizons. As shown in Table 13, inference runtime remains stable (0.014s $\to$ 0.016s) as sequence length grows from 96 to 2048. Memory usage scales linearly with input length.
Table 13: Input length scalability of CTF. We report training runtime (s/iter), peak memory usage (MB), and inference latency as input length grows. Inference runtime remains stable (0.014s $\to$ 0.016s), while memory usage scales linearly with length.
Metric / Input Length 96 512 1024 2048
Training Runtime (s/iter) 0.028 0.048 0.074 0.147
Training Memory (MB) 126.1 403.3 945.8 2921.4
Inference Runtime (s/iter) 0.0140 0.0147 0.0147 0.0159
Inference Memory (MB) 126.1 404.8 945.8 2921.4
Summary.
These results demonstrate that the masking-based attention design in CTF can handle high-density multivariate forecasting and long-horizon inputs with stable inference latency. While memory grows linearly and limits appear at very high channel counts, the model remains efficient in practical regimes.
C.4 Comparison with Modified TimeXer for Asynchronous Sampling
For TimeXer (Wang et al., 2024e), we incorporated channel-wise asynchronously sampled inputs, along with their corresponding sampling information, as exogenous inputs, following the same strategy (shared tokenizers) used in our model. To ensure consistency and reduce redundancy, channels with identical sampling periods shared the same modeling configuration when each was used as the forecasting target. We conducted experiments on the ETT1, ETT2, and SolarWind datasets for Case 1 (Section 5.1), and on the SolarWind dataset for Case 2 (Section 5.1). Results are reported as CMSE and CMAE averaged over four prediction horizons. As summarized in Table 14, our CTF achieves consistently better or comparable performance compared to the modified TimeXer variants.
Table 14: Comparison of CTF with modified TimeXer under asynchronous sampling. Results are reported in CMSE and CMAE averaged over four prediction horizons.
Model CTF (ours) TimeXer (mod.) TimeXer
Dataset / Metric CMSE CMAE CMSE CMAE CMSE CMAE
Case 1 ETT1 0.398 0.411 0.404 0.416 0.420 0.424
ETT2 0.377 0.383 0.374 0.382 0.381 0.389
SolarWind 0.400 0.448 0.417 0.450 0.420 0.468
Case 2 SolarWind 0.440 0.496 0.474 0.504 0.462 0.504
@@ Appendix D Forecasting Results in a Conventional Multivariate Setting
To demonstrate that our model remains effective even under standard multivariate cases, we additionally evaluate it on regularly sampled multivariate time series without any missing values. By aligning the sampling period and patch length across all channels, our model can be directly applied to this conventional setting. This evaluation shows that our model remains applicable beyond the proposed practical settings. For comparison, baseline results (Chen et al., 2025; Wang et al., 2024e; a; Liu et al., 2024c; Huang et al., 2023; Li et al., 2023b; Nie et al., 2023; Zhang & Yan, 2023; Das et al., 2023; Wu et al., 2023; Zeng et al., 2023) are obtained either from the original publications or by reproducing them using the TSLib codebase (Wang et al., 2024d). We compare forecasting performance on five widely used benchmark datasets (Zhou et al., 2021; Wu et al., 2021): ETTh1, ETTh2, ETTm1, ETTm2, and Weather. As shown in Table 15, our model achieves performance comparable to or better than state-of-the-art baselines across all benchmark datasets in the conventional forecasting setting.
@@ Appendix E Full Results of Our Proposed Practical Setting
Table 16 shows the results for Case 1 using asynchronous practical datasets, where datasets with channel-wise asynchronous sampling were linearly interpolated onto a regular sampling grid for the regular time series baselines. Table 17 presents the results for Case 2, which adds block-wise test-time missing inputs on top of asynchronous sampling. All results include error bars based on 5 random seeds to reflect performance variability. Across both cases, our model consistently demonstrates stronger performance than baseline methods.